OpenStudy (anonymous):

5x + 8(x + 5) = 6(x + 2) X=

OpenStudy (anonymous):

hi there! so the strategy here is to multiply everything out, then collect all the X terms on one side, the constants on the other. As a first step, let's try multiplying out. What do you get when you multiply 8(x+5)?

OpenStudy (anonymous):

x = -4

OpenStudy (anonymous):

become a fan please =)!

OpenStudy (anonymous):

what did the person learn just now, GoLearnMath?

OpenStudy (anonymous):

thanks alot

OpenStudy (lilg132):

5x + 8(x + 5) = 6(x + 2) 5x + 8x + 40 = 6x + 12 13x + 40 = 6x + 12 rearrange and bring all x's to one side and numbers to the other side of the '=' sign so you will get 13x - 6x = 12 - 40 7x = -28 now you divide both sides by 7 to get x by it self 7x / 7 = -28 / 7 x = 4

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

x is not 4 it is negative 4 when you multiply 4 with 7 it is 28 not negative 28

OpenStudy (anonymous):

dont forget to click good answer!

OpenStudy (lilg132):

your write my mistake x = -4

OpenStudy (anonymous):

its ok happens all the time

OpenStudy (anonymous):

=)